OpenStudy (anonymous):

use long division to find the quotient and the remanider when P(x)=4x^3-2x^3+x-3 is divided by (x-3)

OpenStudy (anonymous):

\[ 4 x^3-2 x^2+x-3=(x-3) \left(4 x^2+10 x+31\right)+90 \]